# Web-CogReasoner: Towards Knowledge-Induced Cognitive Reasoning for Web Agents

### v1

@@ Abstract
Multimodal large-scale models have significantly advanced the development of web agents, enabling them to perceive and interact with the digital environment in a manner analogous to human cognition. In this paper, we argue that web agents must first acquire sufficient knowledge to effectively engage in cognitive reasoning [^1] Drawing inspiration from Bloom’s educational philosophy, a cornerstone of modern pedagogy.. Therefore, we decompose a web agent’s capabilities into two essential stages: knowledge content learning and cognitive processes. To formalize this, we propose Web-CogKnowledge Framework, which categorizes knowledge into Factual, Conceptual, and Procedural domains. In this framework, knowledge content learning corresponds to the agent’s processes of Memorizing and Understanding, which rely on the former two types knowledge respectively, representing the "what" of learning. Conversely, cognitive processes correspond to Exploring, grounded in Procedural knowledge, defining the "how" of reasoning and action. To facilitate knowledge acquisition, we construct the Web-CogDataset, a structured resource curated from 14 real-world websites, designed to systematically instill the core knowledge necessary for web agent. This dataset serves as the agent’s conceptual grounding—the "nouns" upon which comprehension is built—as well as the basis for learning how to reason and act. Building on this foundation, we operationalize these processes through a novel knowledge-driven Chain-of-Thought (CoT) reasoning framework, developing and training our proposed agent, the Web-CogReasoner. Extensive experimentation reveals its significant superiority over existing models, particularly in its capacity for generalization to unseen tasks where its structured knowledge proves decisive. To facilitate rigorous and systematic evaluation, we introduce the Web-CogBench, a comprehensive evaluation suite designed to assess and compare agent performance across the delineated knowledge domains and cognitive capabilities. Our code and data is open sourced at https://github.com/Gnonymous/Web-CogReasoner.
@@ 1 Introduction
The advent of large-scale models marks a landmark advancement in artificial intelligence. The development of Large Multimodal Models (LMMs), in particular, has vastly expanded the horizons for AI applications. In recent years, AI agents have emerged as a primary vehicle for deploying these models, significantly extending their capabilities. Much like humans, these agents can leverage a diverse array of tools to enhance their skills, leading to remarkable achievements in domains such as code generation hui2024qwen2 ; jiang2024survey , image huang2025diffusion ; bie2024renaissance and video synthesis  assran2025v ; liu2024sora , and academic research. Notably, the application of web agents to surfing the internet has also seen substantial progress.
@@
The evolution of web agents has transitioned from early rule-based systems to modern approaches that widely leverage Large Language Models (LLMs) or Languge Vision Models (LVMs) wang2024survey ; ning2025survey ; zhang2024large ; sapkota2025ai . These LLM-powered agents typically process HTML or Accessibility Tree inputs, formulating tasks and observations into natural language prompts for the model to reason upon and decide the next action. With the integration of LVMs, web agents have gained perceptual abilities akin to human vision, enabling them to comprehend and reason about multimodal content on web pages. As a whole, web agents can be broadly categorized into three types: (1) Text-only agents zhou2023webarena ; li2023camel , which suffer from information loss due to their inability to capture visual cues; (2) Vision-only agents qin2025ui , which rely solely on visual information and thus lose critical structured data; and (3) Hybrid agents koh2024visualwebarena ; he2024openwebvoyager , which utilize both visual and textual information.
LLMs and LMMs are pre-trained on vast amounts of general-domain knowledge, which is undeniably crucial for their foundational capabilities. However, as with any specialized field, this general knowledge is often insufficient, creating a performance bottleneck that caps the potential of the three aforementioned model types. While many existing works have attempted knowledge enhancement or injection, these methods often lack a systematic or theoretically grounded approach. To address this, we draw inspiration from a highly successful paradigm in human education: Bloom’s Taxonomy ormell1974bloom ; conklin2005taxonomy . This framework suggests that effective learning can be deconstructed into two key stages: knowledge content learning and cognitive processes. Accordingly, we propose that training a web agent should follow a similar two-stage methodology. The first stage is dedicated to acquiring knowledge of varying difficulties. The second stage then focuses on the analysis, application, and ultimately, the creative synthesis of that knowledge. This mirrors the human learning trajectory: we first accumulate knowledge through education (Stage 1), and then based on that foundation of knowledge and experience, we learn to apply, innovate, and create (Stage 2).
Therefore, in this paper, we introduce a training paradigm for our model grounded in Bloom’s Taxonomy, bifurcated into two sequential stages: Knowledge Content Learning and Cognitive Processes. The initial stage, Knowledge Content Learning, is dedicated to building a multi-layered knowledge foundation. This foundation consists of three distinct types: Factual knowledge, which includes basic, foundational concepts; Conceptual knowledge, which maps the interrelationships between these facts; and Procedural knowledge, which represents the logical frameworks needed to synthesize factual and conceptual understanding for effective reasoning. To curate this knowledge, we construct Web-CogDataset, collecting metadata from 14 prominent websites and engineered 12 granular tasks designed to systematically instill each knowledge type. In essence, this stage teaches the agent foundational web knowledge and how to apply it to familiar tasks. The subsequent stage, Cognitive Processes, focuses on cultivating the agent’s self-exploring capability. This advanced capability empowers the agent to build upon its learned foundation, enabling it to "learn how to learn" by creatively exploring its own "self-knowledge" to tackle previously unencountered tasks and challenges. To operationalize this, we developed knowledge-guided reasoning templates and utilized imitation learning to train the model. This process explicitly teaches the agent the cognitive faculties required for autonomous learning and creative problem-solving in novel scenarios. To validate our approach, we conducted rigorous evaluations on a suite of general-purpose benchmarks as well as our own. Our findings reveal that the agent trained with our methodology consistently and significantly outperforms state-of-the-art baselines. Crucially, and in alignment with our hypothesis, the performance gap widens dramatically on knowledge-intensive tasks. This demonstrates that, analogous to human learning, the structured acquisition of knowledge is a key determinant of an agent’s ability to excel in complex, domain-specific scenarios.
@@
    Drawing inspiration from Bloom’s taxonomy and established human educational paradigms, we propose the Web-CogKnowledge Framework — a systematic, two-stage training methodology designed to enhance the cognitive capabilities of web agents. Built upon this framework, we develop Web-CogReasoner. Rigorous benchmarking demonstrates that agents trained under our framework achieve a significant performance improvement over current state-of-the-art models.
@@
    We construct the Web-CogDataset, a structured curriculum consisting of 12 fine-grained and progressively challenging tasks. These tasks are meticulously designed to incrementally build the agent’s web knowledge, cognition capability, and higher-order reasoning.
@@ 2 Related Work
Web Agent
Early studies on web page understanding have concentrated on leveraging structured text. The work gur2022understanding defined crucial HTML understanding tasks, such as semantic classification, description generation, and autonomous web navigation, and trained models to comprehend the structure and content of HTML. AutoWebGLMlai2024autowebglm introduced an HTML simplification algorithm. Through curriculum learning, it first trained the model using collected web page data to recognize web page structures and understand the functions of different interactive components, like text boxes and buttons. Subsequently, the model learned to execute simple web tasks and decompose complex ones into subtasks, enabling the complete execution of web tasks from simple to complex vision-based Web Page Understanding.
Recent research has explored the use of visual information for web page understanding. SeeClickcheng2024seeclick designed specialized pre-training tasks, including text-to-point and text-to-bounding box localization tasks, as well as point-to-text and bbox-to-text generation tasks. These tasks facilitated the model’s learning of the correlation between the positions of web elements and their textual descriptions, enhancing the model’s ability to localize and understand web elements. CogAgenthong2024cogagent introduced a high-resolution cross-module, enabling the model to recognize tiny elements and text in GUIs, and constructed a large-scale GUI annotated dataset for GUI grounding tasks, achieving excellent performance in multiple visual question answering (VQA) benchmarks and GUI navigation task datasets. OmniParserwan2024omniparser integrated three typical visually-situated text parsing tasks—text spotting, key information extraction, and table recognition—within a unified framework, providing a more efficient solution for GUI Agents to handle complex tasks. Unlike many existing agent frameworks that rely on heavily wrapped commercial models (e.g., GPT-4o) with expert-crafted prompts and workflows, UI-TARSqin2025ui adopted an end-to-end architectural design, directly perceiving screenshots, executing reasoning processes, and autonomously generating valid actions. UGroundgou2024navigating developed a universal visual grounding model using a GUI dataset with 10 million elements, demonstrating outstanding performance in desktop tasks and mobile tasks.
Multimodal approaches have emerged as a powerful way to comprehensively understand web pages. WebVoyagerhe2024webvoyager utilized the GPT-4V-ACT tool to overlay bounding boxes and numerical labels of interactive elements on web page screenshots. Additionally, it complemented web page visual information with the Accessibility Tree, helping the model gain a more comprehensive understanding of web page content. SeeActzheng2024gpt directly harnessed the powerful visual and language understanding capabilities of GPT-4V and explored various grounding strategies to map the text plans generated by the model into specific web operations. Tuning-free Attention-driven Grounding (TAG) xu2025attention method further accomplishes this task without the need for additional fine-tuning by leveraging the inherent attention patterns in pretrained MLLMs.
Web Agent Evaluation
To evaluate GUI agents’capabilities in web interactions, benchmarks are broadly categorized into web browsing tasks and web understanding assessments. For web browsing, offline benchmarks like Mind2Webdeng2023mind2web , Multimodal-Mind2Web, AutoWebBenchlai2024autowebglm and WebVLN-v1chen2024webvln evaluate models’ability to perform multi-step web tasks (e.g., booking flights, filling dynamic forms) using static datasets that include webpage content along with corresponding interaction trajectories. In contrast to offline benchmarks, online environments such as Mini-WoB++liu2018reinforcement , Webshopyao2022webshop , and WebArenazhou2023webarena enable real-time interaction testing. Mini-WoB++liu2018reinforcement focuses on low-level UI operations (e.g., clicking buttons, handling pop-ups) within simplified webpages. By comparison, Webshopyao2022webshop and WebArenazhou2023webarena replicate complex, real-world tasks—such as e-commerce transactions and Wikipedia editing—to evaluate task completion performance. Building on this, VisualWebArenakoh2024visualwebarena incorporates multimodal inputs to assess models’ ability to handle vision-augmented interactions in dynamic web environments. To evaluate deeper web understanding, benchmarks are designed to target specific cognitive abilities. WEBQAchang2022webqa assesses a model’s capacity for open-domain, multi-hop reasoning. In contrast, ScreenQAhsiao2022screenqa and the ScreenAIbaechler2024screenai suite focus on screen-centric comprehension. ScreenQAhsiao2022screenqa evaluates a model’s ability to recognize UI elements and interpret context by answering questions based on screenshots. ScreenAIhsiao2022screenqa further extends this with three progressively challenging subtasks: Screen Annotation (SA), which involves spatial layout tagging; ScreenQA Short, which tests concise fact extraction; and Complex ScreenQA, which requires multi-step reasoning across diverse screen layouts. Together, these tasks assess fine-grained layout understanding, semantic interpretation, and practical problem-solving in visually dense interfaces.
@@ 3.1 Bloom’s Taxonomy
The Bloom’s Taxonomy  [^2] https://fctl.ucf.edu/teaching-resources/course-design/blooms-taxonomy/. (Anderson & Krathwohl, 2001) presents a two-dimensional pedagogical design, i.e., knowledge content learning and cognitive process dimension, which demonstrates a "shallow-to-deep" instructional methodology. This approach posits that effective learning is not monolithic but is a structured progression through distinct types of knowledge, beginning with foundational elements and culminating in high-level self-awareness. This educational strategy ensures that learners first establish a solid base of facts and concepts before advancing to the execution of complex procedures and, ultimately, the application of reflective metacognitive strategies.
This pedagogical progression from simple to complex can be formally outlined through four hierarchical types of knowledge: Factual Knowledge, Conceptual Knowledge, Procedural Knowledge and Metacognitive Knowledge. More details can be found in Appendix A.1.1.
@@ 3.2 Web-CogKnowledge
    Factual knowledge: concrete information extracted from web content. Such as identifying the attributes of individual web elements and predicting the immediate, direct consequences of a single interaction;
@@
    Conceptual knowledge: semantic relationships and abstract patterns underlying webpage content and structure. Such as inferring the function of interface components, comprehending the overall purpose and structure of a webpage, and interpreting its multimodal content;
@@
    Procedural knowledge: actionable know-how for accomplishing specific tasks through interaction, including planning, decision-making, and sequential execution. Such as executing goal-oriented action sequences, inferring user intent from observed behaviors, and handling unexpected interruptions to complete complex tasks;
@@
Figure 2: Statistics of Selected Websites by Category.
To fully leverage the potential of Web-CogKnowledge, leading to the ultimate realization of the Web-CogKnowledge Framework. We first collect multimodal metadata from 14 real-world websites (see Figure 2) and then design Web-CogDataset with diverse set of web tasks for Web-Agent training in Section 3.3. And finally construct Web-CogBench in Section 3.4, built from meticulously chosen subsets. Specifically, we curate the metacognitive knowledge and Web-CogReasoner learning and training process in Section 4 due to its importance and complexity. More details about data collection and cleaning process can be found in Appendix A.1.3.
@@ 3.3 Web-CogDataset
By selectively crawling metadata from the aforementioned websites and leveraging the hierarchical structure of Web-CogKnowledge, we define a series of tasks corresponding to different levels of knowledge granularity. This process results in the construction of a comprehensive and hierarchically structured dataset, referred to as Web-CogDataset. The task definitions, organized according to the knowledge hierarchy, are introduced below. Detailed statistics and illustrative examples for Web-Dataset are provided in Table 1 and Figure 3. For additional implementation details, please refer to the Appendix A.2.
Table 1: Statistics of the Web-CogDataset.
Knowledge Task Statistics
Subtotal Total
Factual Web Knowledge Element Attribute Recognition 37K 81K
Sub-elements Prediction 1K
Page Change Prediction 24K
Next Page Prediction 18K
Source Element Prediction 1K
Conceptual Web Knowledge Element Understanding 40K 62K
WebPage Understanding 7K
Caption & QA 15K
Procedural Web Knowledge User’s Intention Prediction 2K 27K
Popup Close 1K
Single-Step Web Task 17K
Noisy Multi-Steps Web Task 7K
Factual Web Knowledge Tasks Overview
The Factual Web Knowledge tasks are designed to establish the agent’s core perceptual abilities by operationalizing our definition of factual knowledge. These five tasks train the agent to identify the attributes of individual web elements and predict the immediate, direct consequences of a single interaction. By focusing on extracting concrete, visually grounded information—rather than abstract reasoning—this task category ensures the agent builds a foundational understanding of the web’s basic cause-and-effect mechanics. This layer of knowledge is the essential prerequisite for any subsequent conceptual or procedural reasoning.
  • Element Attribute Recognition: Given a screenshot with a highlighted interactive element, the model predicts its semantic role (e.g., button, link) and accessible name (e.g., "Submit", "Search"), relying solely on visual cues.
  • Sub-elements Prediction: The model infers the dynamic hierarchical structure of a web interface by predicting the set of sub-elements that appear upon interacting with a highlighted parent element.
  • Page Change Prediction: The model anticipates the visual changes that occur after clicking on a specific element, formulating an open-ended response that describes the expected dynamic behavior of the page.
  • Next Page Prediction: The model predicts the subsequent page that results from interacting with a specific element on the current page. To enhance generalization, we designed two types of tasks: multiple-choice questions and open-ended responses.
  • Source Element Prediction: Given two screenshots—the current page and the resulting target page. Then the model identifies which of the visually marked elements on the current page leads to the target, simulating visual cause-and-effect reasoning.
Figure 3: Illustration of Web-CogDataset.
Conceptual Web Knowledge Tasks Overview
Building upon the factual foundation, the Conceptual Web Knowledge tasks are designed to move the agent from perception to comprehension. These tasks train the agent to synthesize discrete facts into meaningful patterns and understand the semantic relationships underlying a webpage’s content and structure. Rather than merely identifying elements, the agent learns to infer their functions, comprehend the webpage’s overall purpose and organization, and interpret its multimodal content. This layer of conceptual understanding is critical, as it enables the agent to grasp the "why" behind the interface, transforming it from a collection of static elements into a coherent, interactive system.
  • Caption & QA: The model generates captions and answers questions about embedded images and WebPages, assessing its visual reasoning and holistic understanding of multimodal webpage information (e.g., layout, text). Contributes to further enhancing the model’s understanding of WebPages, while preserving its open conversational abilities with web content. The orignal data is drawn from the Multi-UI liu2024harnessing dataset.
Procedural Web Knowledge Tasks Overview
With both factual and conceptual knowledge established, the Procedural Web Knowledge tasks are designed to transition the agent from understanding to action. These tasks train the agent to formulate and execute goal-oriented plans by leveraging its actionable know-how for accomplishing specific tasks through interaction. The agent learns to decompose complex user goals into sequential steps, handle unexpected interruptions like popups, and orchestrate a series of interactions to achieve a desired outcome. This final knowledge layer is paramount, as it empowers the agent to apply its understanding strategically, converting passive knowledge into active, successful task completion.
  • User’s Intention Prediction: The model infers high-level user intent from a sequence of webpage screenshots representing an interaction trajectory, requiring visual understanding and temporal reasoning. The task is built on the MultiModal-Mind2Web deng2023mind2web dataset, mapping screenshot sequences to natural language instructions.
  • Popup Close: The model identifies and dismisses popups (e.g., notification modals, login forms) on synthesized webpage screenshots, using a dataset of 51 popup components from JS Design  [^3] https://js.design/. overlaid on OpenWebVoyager he2024openwebvoyager webpages, with combinatorial augmentation of closing strategies for diverse training.
  • Single-Step Web Task: The model executes goal-directed web interactions to perform atomic web actions, using annotated screenshots from the MultiUI liu2024harnessing dataset, with tasks validated for accuracy.
  • Noisy Multi-Steps Web Task: The model perform complex, goal-oriented interactions across multiple steps, based on unified OpenWebVoyager he2024openwebvoyager trajectories, with a standardized action space for consistent evaluation. To test robustness, some tasks include UI interruptions (e.g., popups) that must be dismissed before resuming the task.
Table 2: Detailed statistics of Web-CogBench.

### v2

@@ Abstract
Multimodal large-scale models have significantly advanced the development of web agents, enabling them to perceive and interact with the digital environment in a manner analogous to human cognition. In this paper, we argue that web agents must first acquire sufficient knowledge to effectively engage in cognitive reasoning [^1] Drawing inspiration from Bloom’s educational philosophy, a cornerstone of modern pedagogy.. Therefore, we decompose a web agent’s capabilities into two essential phases: knowledge content learning and cognitive processes. To formalize this, we propose the Web-CogKnowledge Framework, which categorizes knowledge into Factual, Conceptual, and Procedural domains. In this framework, knowledge content learning corresponds to the agent’s processes of Memorizing and Understanding, which rely on the first two types of knowledge respectively, representing the ”what” of learning. Conversely, cognitive processes correspond to Exploring, grounded in Procedural knowledge, defining the ”how” of reasoning and action. To facilitate knowledge acquisition, we construct Web-CogDataset, a structured resource curated from 14 real-world websites, designed to systematically instill the core knowledge necessary for a web agent. This dataset serves as the agent’s conceptual grounding—the ”nouns” upon which comprehension is built—as well as the basis for learning how to reason and act. Building on this foundation, we operationalize these processes through a novel knowledge-driven Chain-of-Thought (CoT) reasoning framework, developing and training our proposed agent, Web-CogReasoner. Extensive experimentation reveals its significant superiority over existing models, particularly in its capacity for generalization to unseen tasks where its structured knowledge proves decisive. To facilitate rigorous and systematic evaluation, we introduce Web-CogBench, a comprehensive evaluation suite designed to assess and compare agent performance across the delineated knowledge domains and cognitive capabilities. Our code and data is open sourced at https://github.com/Gnonymous/Web-CogReasoner.
@@ 1 Introduction
The advent of large-scale models marks a milestone in artificial intelligence, with Large Multimodal Models (LMMs) greatly expanding application horizons. AI agents have become the primary vehicle for deploying these models, enabling capabilities in code generation (Hui et al., 2024; Jiang et al., 2024), image and video synthesis (Huang et al., 2025; Bie et al., 2024; Assran et al., 2025; Liu et al., 2024c), and academic research (Li et al., 2025; Zhang et al., 2025). Recent progress has also highlighted the growing importance of web agents.
@@
Web agents have evolved from early rule-based systems to modern approaches leveraging Large Language Models (LLMs) and Language Vision Models (LVMs) (Wang et al., 2024; Ning et al., 2025; Zhang et al., 2024; Sapkota et al., 2025). LLM-powered agents typically convert HTML or Accessibility Tree inputs into natural language prompts for reasoning and action. With LVMs, agents gain perceptual abilities akin to human vision, allowing them to process multimodal content on web pages. Broadly, web agents can be categorized as: (1) text-only (Zhou et al., 2023; Li et al., 2023), which miss visual cues; (2) vision-only (Qin et al., 2025), which lack structured data; and (3) hybrid (Koh et al., 2024; He et al., 2024b), which integrate both modalities.
LLMs and LVMs pre-trained on general-domain knowledge provide strong foundations but remain limited in specialized tasks, creating performance bottlenecks. Prior knowledge-enhancement methods often lack systematic or theoretical grounding. To address this, we draw inspiration from Bloom’s Taxonomy (Ormell, 1974; Conklin, 2005), which divides learning into two phases: Knowledge Content Learning and Cognitive Processes.
In our paradigm, the first phase builds a multi-layered foundation: Factual Knowledge, covering basic concepts, and Conceptual Knowledge, capturing their interrelations. This equips the agent with core web knowledge and its application to familiar tasks. The second phase develops Procedural Knowledge, providing logical frameworks to synthesize prior knowledge for reasoning and exploration. This enables the agent to “learn how to learn,” creatively leveraging self-knowledge to solve novel challenges. This mirrors the human learning trajectory: we first accumulate knowledge through education (Phase 1), and then based on that foundation of knowledge and experience, we learn to apply, innovate, and create (Phase 2).
To support this, we construct Web-CogDataset from 14 prominent websites with 12 fine-grained tasks, and design knowledge-guided reasoning templates combined with imitation learning to instill the required cognitive faculties. Rigorous evaluations on public and in-domain benchmarks show that our method consistently surpasses state-of-the-art baselines, with the performance advantage especially pronounced in knowledge-intensive tasks. These results confirm that structured knowledge acquisition is crucial for enabling agents to excel in complex, domain-specific scenarios.
@@
    Drawing inspiration from Bloom’s taxonomy and established human educational paradigms, we propose the Web-CogKnowledge Framework, a systematic, two-phase training methodology designed to enhance the cognitive capabilities of web agents. As shown in Figure 1, built upon this framework, we develop Web-CogReasoner. Rigorous benchmarking demonstrates that agents trained under our framework achieve a significant performance improvement over current state-of-the-art models.
@@
    We construct Web-CogDataset, a structured curriculum consisting of 12 fine-grained and progressively challenging tasks. These tasks are meticulously designed to incrementally build the agent’s web knowledge, cognition capability, and higher-order reasoning.
@@ 2 Related Work
2.1 Web Agent
Early work on web understanding focused on structured HTML, addressing tasks like semantic classification, description generation, and navigation (Gur et al., 2022), with AutoWebGLM (Lai et al., 2024) further applying curriculum learning for structure recognition, component understanding, and progressively complex task execution. More recent studies leverage visual signals: SeeClick (Cheng et al., 2024) linked elements to textual descriptions to enhance localization, CogAgent (Hong et al., 2024) combined high-resolution cross-module modeling with a large GUI dataset for VQA and navigation, OmniParser (Wan et al., 2024) unified text spotting, extraction, and table recognition, UI-TARS (Qin et al., 2025) directly maps screenshots to actions, and UGround (Gou et al., 2024) trained on 10M GUI elements for robust desktop and mobile performance. Multimodal approaches integrate text and vision: WebVoyager (He et al., 2024a) combines screenshots with bounding boxes and accessibility trees, SeeAct (Zheng et al., 2024) grounds text plans via GPT-4V, and TAG (Xu et al., 2025) exploits pretrained attention for grounding without fine-tuning.
2.2 Web Agent Evaluation
Benchmarks are categorized into browsing and understanding. For browsing, offline datasets like Mind2Web (Deng et al., 2023), Multimodal-Mind2Web, AutoWebBench (Lai et al., 2024), and WebVLN-v1 (Chen et al., 2024) test multi-step task execution, while online environments such as Mini-WoB++ (Liu et al., 2018), Webshop (Yao et al., 2022), and WebArena (Zhou et al., 2023) allow real-time evaluation. Mini-WoB++ emphasizes low-level UI operations, whereas Webshop and WebArena simulate complex tasks. VisualWebArena (Koh et al., 2024) further adds multimodal inputs for dynamic interactions.
For web understanding, WEBQA (Chang et al., 2022) evaluates open-domain multi-hop reasoning. ScreenQA (Hsiao et al., 2022) and ScreenAI (Baechler et al., 2024) focus on screen comprehension, with ScreenQA targeting UI recognition and contextual QA, while ScreenAI extends this into three subtasks: Screen Annotation, ScreenQA Short, and Complex ScreenQA. Together, they assess layout understanding, semantic interpretation, and reasoning in visually dense interfaces.
@@ 3.1 Bloom’s Taxonomy
The Bloom’s Taxonomy [^2] https://fctl.ucf.edu/teaching-resources/course-design/blooms-taxonomy/ (Anderson & Krathwohl, 2001) presents a two-dimensional framework: knowledge content learning and cognitive processes, embodying a ”shallow-to-deep” instructional methodology. It structures learning from foundational facts and concepts to complex procedures, ensuring a solid knowledge base before higher-level reasoning.
This progression can be formalized through four hierarchical types of knowledge: Factual, Conceptual, and Procedural Knowledge. Further details are in Appendix A.1.1.
Figure 2: Illustration of Web-CogDataset.
@@ 3.2 Web-CogKnowledge
    Factual knowledge: concrete information extracted from web contents, such as identifying the attributes of individual web elements and predicting the immediate, direct consequences of a single interaction.
@@
    Conceptual knowledge: semantic relationships and abstract patterns underlying webpage content and structures, such as inferring the function of interface components, comprehending the overall purpose and structure of a webpage, and interpreting its multimodal content.
@@
    Procedural knowledge: actionable know-how for accomplishing specific tasks through interaction, including planning, decision-making, and sequential execution. Such as executing goal-oriented action sequences, inferring user intent from observed behaviors, and handling unexpected interruptions to complete complex tasks.
@@
To fully leverage the potential of Web-CogKnowledge, leading to the ultimate realization of the Web-CogKnowledge Framework, we first collect multimodal metadata from 14 real-world websites (see Figure 5) and then design Web-CogDataset with diverse sets of web tasks for web-agent training in Section 3.3, and finally construct Web-CogBench in Section 3.4, built from meticulously chosen subsets. Specifically, we curate Web-CogReasoner learning and training process in Section 4 due to its importance and complexity. More details about data collection and cleaning process can be found in Appendix A.1.5.
@@ 3.3 Web-CogDataset
By selectively crawling metadata from 14 representative websites and aligning it with the hierarchical design of Web-CogKnowledge, we construct Web-CogDataset, a large-scale, hierarchically structured dataset tailored for knowledge-centric web reasoning. The dataset spans three layers of knowledge: Factual, Conceptual, and Procedural. Each mapped to carefully designed task families that progressively cultivate perception, comprehension, and action-oriented reasoning.
As illustrated in Figure 2, these tasks together form a coherent pipeline that transitions agents from identifying elemental attributes, to grasping semantic patterns and page structures, and finally to executing complex, goal-directed interactions under realistic constraints. This organization mirrors human learning trajectories, ensuring that higher-order reasoning is built on solid perceptual and conceptual foundations.
Detailed task definitions, implementation protocols, and examples are provided in Appendix A.2. Dataset statistics are reported in Table 13. To ensure data quality and diversity, we further analyze the dataset composition and annotation reliability in Appendix A.1.4.
3.4 Web-CogBench
To evaluate the cognitive capabilities enabled by our knowledge-centric framework, we introduce Web-CogBench. While our training dataset is organized by knowledge type (Factual, Conceptual, Procedural), Web-CogBench measures agent performance across three corresponding abilities: Memorizing, Understanding, and Exploring. Curated from a representative subset of Web-CogDataset, it assesses how effectively an agent applies learned knowledge in complex web contexts. Detailed statistics are in Table 1, and the evaluation dimensions align with our hierarchical knowledge framework. A complete definition of all tasks is provided in the appendix A.2.2.
Table 1: Statistics of Web-CogBench.